# The impact of stellar bars on quenching star formation: insights from a spatially resolved analysis in the local Universe

### v1

@@ Abstract
Context. Stellar bars are common morphological structures in the local Universe: according to optical and NIR surveys, they are present in about two-thirds of disc galaxies. These elongated structures are also believed to play a crucial role in secular evolutionary processes, since they are able to efficiently redistribute gas, stars and angular momentum within their hosts, although it remains unclear whether they enhance or suppress star formation. A useful tool to investigate such an ambiguity is the Main Sequence (MS) relation, which tightly links stellar mass () and star formation rate (SFR).
Aims. The main goal of this work is to explore star formation processes in barred galaxies, in order to assess the relevance of possible bar quenching effects on the typical log-linear trend of the resolved MS.
Methods. To this purpose, we carry out a spatially resolved analysis on sub-kpc scales for a sample of six nearby barred galaxies. Multi-wavelength photometric data (from far-UV to far-IR) are collected from the DustPedia database and a panchromatic Spectral Energy Distribution (SED) fitting procedure is applied on square apertures of fixed angular size ( ), making use of the magphys code.
Results. For each galaxy we obtain the distributions of stellar mass and star formation rate surface densities and relate them in the – plane deriving the spatially resolved MS relation. Although significant galaxy-to-galaxy variations are in place, we infer the presence of a common anti-correlation track in correspondence with the bar-hosting region, which shows systematically lower values of SFR. Such a central quiescent signature can be interpreted as the result of a bar-driven depletion of gas reservoirs and a consequent halt of star formation. This seems to point in the direction of an inside-out quenching scenario.
@@
Internal processes typically dominate the late-stage evolution of galaxies, when gravitational interactions and mergers become less frequent as the Universe expands (Kormendy & Kennicutt, 2004). Stellar bars are among the main internal drivers of the secular evolution of disc galaxies, since they are able to shape the current properties of their hosts redistributing gas, stars and angular momentum within them (e.g., Athanassoula & Misiriotis, 2002; Sheth et al., 2005; Debattista et al., 2006; Athanassoula, 2013; Di Matteo et al., 2013; Sellwood, 2014).
@@
The pivotal role of bars in improving our understanding of galaxy evolution is also supported by the fact that they are common morphological structures in the local Universe. According to optical and near-infrared (NIR) surveys, between and of nearby disc galaxies host stellar bars (e.g., Eskridge et al., 2000; Barazza et al., 2008; Aguerri et al., 2009; Nair & Abraham, 2010; Buta et al., 2015; Erwin, 2018). The evolution of this fraction over cosmic time may shed light on the epoch of bar formation. In particular, as found by several observational studies, there seems to be a consensus on the decline of the barred galaxy fraction from the local Universe up to by at least a factor of two (e.g., Sheth et al., 2008; Melvin et al., 2014). More recently, the unprecedented sensitivity and spatial resolution in the NIR of the James Webb Space Telescope (JWST) have given us the chance to infer the evolution of the bar fraction at higher redshifts. Prominent elongated structures have been identified in disc galaxies up to (e.g., Guo et al., 2023; Costantin et al., 2023; Amvrosiadis et al., 2024), while Le Conte et al. (2023) presented a full census of observable stellar bars at . They pointed out that the incidence of bars in disc galaxies decreases from to going from to . Consistent results were also obtained in a fully cosmological context, thanks to a new generation of high-resolution simulations (e.g., Fragkoudi et al., 2020; Zhao et al., 2020; Rosas-Guevara et al., 2022).
@@
The fact that both observational studies and simulations present evidence of a significant population of barred galaxies up to suggests that stellar bars are already in place in the early Universe, thus hinting at the possibility that bar-driven secular processes could potentially influence galaxy evolution over billions of years. Bars are thought to form over time scales of the order of a hundred million years, once a dynamically cold and rotationally supported disc has settled (e.g., Ostriker & Peebles, 1973; Sellwood & Wilkinson, 1993). The onset of bar formation occurs mainly when the stellar orbits of massive disc galaxies deviate from a circular path due to instabilities.
@@
Whether the growth of stellar bars enhance or suppress star formation is still debated (e.g., Kim et al., 2017; Wang et al., 2020). However, there are indications supporting the hypothesis that bars can suppress star formation (bar quenching) in a region spanning from the central sub-kiloparsec (sub-kpc) to the end of the bar itself (e.g., Masters et al., 2010, 2012; Gavazzi et al., 2015; James & Percival, 2016; George et al., 2019a; Rosas-Guevara et al., 2020; Zhang et al., 2021), thus contributing to the observed transition from active star-forming galaxies to passive ones (e.g., Man & Belli, 2018). This could be due to the bar-induced shock and shear, which stabilise the gas against collapse by increasing turbulence and, in turn, resulting in the inhibition of star formation (e.g., Haywood et al., 2016; Khoperskov et al., 2018). Another possible mechanism is related to the redistribution of cold gas along the length of the stellar bar. The torque produced by this morphological structure perturbs the interstellar medium (ISM), driving gas and dust inflow towards the galactic centre (e.g., Casasola et al., 2011; Combes et al., 2013; Spinoso et al., 2017). As a consequence, this boosts the star formation in the inner sub-kpc-scale area and depletes the bar hosting region of fuel for further star formation (e.g., George et al., 2019a, 2020; Newnham et al., 2020).
@@
The bar-quenching scenario, suggested by many recent observations and simulations, predicts that bars can be responsible for the onset of a fast star formation rate phase at their early evolutionary stages, potentially triggering starburst events or fuelling an Active Galactic Nucleus (AGN), accompanied by a subsequent rapid consumption of gas, thus inducing a premature quenching phase in their host galaxies (e.g., Spinoso et al., 2017; Fraser-McKelvie et al., 2020a, b; George & Subramanian, 2021; Géron et al., 2021). This could explain why some literature works agree on the fact that barred galaxies have higher SFR than unbarred ones, especially at the centre, along the length of the bar and within the circumnuclear region (e.g., Sheth et al., 2005; Zurita & Pérez, 2008; Ellison et al., 2011; Coelho & Gadotti, 2011; Prieto et al., 2019; Lin et al., 2020). While many other studies pointed out that bars are most likely to be found in quiescent galaxies, which are typically more massive, redder, and gas-poor (e.g., Vera et al., 2016; Cervantes Sodi, 2017; Géron et al., 2021).
@@
In this context, it is worthwhile to further explore the impact of bar quenching by studying the star formation properties and their dependence on the presence of the bar, especially through the existing empirical correlations between the main structural parameters of galaxies. In particular, the stellar mass () – star formation rate (SFR) relation is one of the most useful tools in modern astrophysics (e.g., Brinchmann et al., 2004; Noeske et al., 2007a, b; Salim et al., 2007). It is usually named the Main Sequence (MS) of star-forming galaxies (SFGs) and it shows that more massive galaxies host larger amounts of star formation than lower-mass ones. Many literature works hint that this log-linear correlation is valid, with an intrinsic scatter of 0.3 dex, for local as well as for distant galaxies (up to z 6), for a wide range of stellar masses and considering different SFR tracers (e.g., Daddi et al., 2007; Pannella et al., 2009; Santini et al., 2009; Oliver et al., 2010; Elbaz et al., 2011; Lee et al., 2012; Whitaker et al., 2012; Moustakas et al., 2013; Whitaker et al., 2014; Tasca et al., 2015; Barro et al., 2017; Santini et al., 2017; Pearson et al., 2018; Popesso et al., 2019b; Daddi et al., 2022; Popesso et al., 2023). Galaxies exhibiting various rates of stellar production (starbursts, SFGs, and passive galaxies) populate different regions of the – SFR plane. Therefore, their position with respect to the MS can be used as a diagnostic tool to identify their state of star formation and to figure out the origin of the colour bimodality of galaxies (e.g., Rodighiero et al., 2011; Renzini & Peng, 2015; Bisigello et al., 2018; Popesso et al., 2019a).
@@
A significant limitation in the understanding of the – SFR relation is that many studies have relied on the integrated properties of galaxies without distinguishing among their morphological components. A more advanced approach involves characterising this correlation on smaller scales, employing spatially resolved data. Thanks to high-spatial-resolution imaging and to the advent of integral field spectroscopy (IFS) facilities, it was discovered that the MS relation is valid also locally, at kpc and sub-kpc scales (e.g., Sánchez et al., 2013; Wuyts et al., 2013; Cano-Díaz et al., 2016; Abdurro’uf & Akiyama, 2017; Hsieh et al., 2017; Lin et al., 2019; Enia et al., 2020; Ellison et al., 2021; Casasola et al., 2022). This implies that the star formation activity originates from physical mechanisms that could be universal across various physical scales. There is a general consensus on the existence of such a ”resolved” Main Sequence relation between the locally measured star formation rate surface density () and the stellar mass surface density (). However, its slope, zero-point and scatter are subject to significant galaxy-to-galaxy variations, depending on the different morphological structures, SFR tracers, dust extinction correction, and fitting procedures (e.g., Abdurro’uf & Akiyama, 2017; Vulcani et al., 2019; Pessa et al., 2021, 2022; Casasola et al., 2022).
@@
With the aim of probing the relevance of bar-driven quenching, several observational works and simulations dealt with the analysis of the integrated MS relation of statistically large samples of barred galaxies (e.g., Vera et al., 2016; George et al., 2019b; Fraser-McKelvie et al., 2020a; George & Subramanian, 2021; Rosas-Guevara et al., 2020, 2022). Instead, as far as we know, there are just a few studies that investigated the spatially resolved MS across different galactic structures, including the bar (Pessa et al., 2021, 2022). Moreover, previous literature works are almost based on spectroscopic data or UV-to-optical tracers of SFR (see Sánchez, 2020, for a review).
@@
In this work, we further investigate the MS relation at sub-kpc scales for six barred galaxies in the local Universe, exploiting the power of multi-wavelength photometric data (from far-UV to far-IR) collected in the DustPedia archive (Davies et al., 2017; Clark et al., 2018) and building on the methodology proposed in Enia et al. (2020). With this approach, we can retrieve a reliable measure of the SFR, taking into account both the unobscured and the dust-obscured components, once a Spectral Energy Distribution (SED) fitting procedure is applied. The final aim of this study is to get a more comprehensive view of the star formation processes in barred galaxies and to understand whether the presence of the bar affects the typical log-linear trend of the resolved MS, at least in the inner regions.
@@
The paper is structured as follows: in Section 2, we briefly describe the data used in this work and our sample selection; in Section 3, we report the main steps of the employed methodology for deriving and SFR; in Section 4, we present our results concerning the spatially resolved MS; in Section 5, we discuss the implications of our findings. Finally, in Section 6, we summarise our work and the main outcomes. Throughout the paper, we will assume a CDM cosmology with parameters , , (Planck Collaboration et al., 2016) and a Chabrier (2003) Initial Mass Function (IMF).
@@ 2.1 DustPedia archive
DustPedia111The DustPedia website is available at http://dustpedia.astro.noa.gr/ is a research project that was developed to enable a comprehensive characterisation of the cosmic dust in the local Universe and of its effects on multi-wavelength observations of ISM in galaxies (Davies et al., 2017).
The DustPedia sample consists of 875 galaxies that were observed by Herschel (Pilbratt et al., 2010) and that lie within 50 Mpc, with an optical diameter D25222D25 is the major axis of the 25-th level isophote, which is the one at which the value of the surface brightness falls below 25mag/arcsec2 in the B band. greater than . The complete data set, which is publicly available to the scientific community, contains standardised multi-wavelength imagery and photometry of the whole sample of galaxies, including both custom reductions and archival data, spanning over five orders of magnitude in wavelength from ultraviolet to microwave (Clark et al., 2018).
@@ 2.2 Sample selection
To properly select a sample of barred spiral galaxies from the DustPedia database, we first put some constraints on the Hubble stage index T444The Hubble stage index T is a numerical value, running from -6 to 10, that is assigned to each morphological class of galaxies in the De Vaucouleurs scheme. Negative values correspond to early-type galaxies (ellipticals and lenticulars), while positive numbers correspond to late-type galaxies (spirals and irregulars)., restricting its range between 2 and 7. In this way we exclude early-type galaxies, intermediate classifications between lenticular and spirals and also the irregular ones. Then we apply a cutoff in distance, at approximately 20 Mpc, and also impose a lower limit for the optical diameter D25, considering only galaxies with D . Both these conditions ensure a selection of sufficiently close and extended galaxies, so that their central regions hosting the bar can be properly investigated with enough resolution. Furthermore, we focus on sources that have a disc inclination of 65°, to avoid projection corrections. Besides all this, we select only spiral galaxies that show a bar structure based on their known morphological classification. We also exclude objects with peculiar structures and those hosting obvious AGN (according to the already existing literature). Indeed, the presence of a strong nuclear activity would largely affect the emission due to the star formation in most of the galaxy and especially in the central region, which is the one to be deeply investigated in order to probe the relevance of the bar-driven quenching. In the end, thanks to all the outlined constraints, we obtain a sample as free as possible from any contamination, enabling us to more effectively distinguish the impact of the bar.
@@
The final sample consists of six objects: NGC 1566, NGC 3351, NGC 3953, NGC 4579, NGC 4725, NGC 5236. Distances, inclinations, sizes and integrated properties of the galaxies under consideration are presented below and summarised in Table 1. All the sources are identified by their own equatorial coordinates (Right Ascension and Declination) at epoch J2000. For each galaxy, along with the morphological type, we report: luminosity distance in Mpc, inclination, R25 (semi-major axis of the 25-th level isophote) size in kpc. All these quantities are taken from the HyperLEDA database (Makarov et al., 2014) and are those adopted by the DustPedia collaboration. Moreover, we report the bar radius Rbar in kpc, as determined in previous studies based on photometric decompositions of images from the Spitzer Survey of Stellar Structure in Galaxies (S4G). In particular, we retrieve the bar radius of NGC 1566 and NGC 3351 from Muñoz-Mateos et al. (2013), while for all the other galaxies we refer to the work of Salo et al. (2015). Then, we show the total stellar mass and the total star formation rate, with the corresponding uncertainties, of all the selected sources. These integrated properties are expressed respectively in units of and of yr-1 and are taken from the results of Nersesian et al. (2019), paying attention to properly rescale them in accordance with a Chabrier (2003) IMF. They derived the global properties of most of the DustPedia sample using the aperture-matched photometry generated by the Comprehensive Adaptable Aperture Photometry Routine (CAAPR), which is presented in Clark et al. (2018). On this photometric dataset, they modelled the panchromatic SED employing the cigale fitting code (Boquien et al., 2019) and assuming the Salpeter (1955) IMF. We also indicate the size in kpc of the square apertures adopted for the spatially resolved analysis presented in this paper (i.e., the size of the cells). In particular, this value is related to the so-called pixel-by-pixel SED fitting procedure that is applied to a grid of square cells with side, which corresponds to different physical sizes depending on the galaxies’ distances. More details about the methodology are explained in Sect. 3.

### v2

# The impact of stellar bars on star-formation quenching: Insights from a spatially resolved analysis in the local Universe

@@ Abstract
Context. Stellar bars are common morphological structures in the local Universe; according to optical and NIR surveys, they are present in about two-thirds of disc galaxies. These elongated structures are also believed to play a crucial role in secular evolutionary processes, because they are able to efficiently redistribute gas, stars, and angular momentum within their hosts, although it remains unclear as to whether they enhance or suppress star formation. A useful tool to investigate this ambiguity is the main sequence (MS) relation, which tightly links stellar mass () and star formation rate (SFR).
Aims. The main goal of this work is to explore star-formation processes in barred galaxies in order to assess the relevance of bars in star-formation quenching and whether or not they affect the typical log-linear trend of the resolved MS.
Methods. To this purpose, we carried out a spatially resolved analysis on subkiloparsec (subkpc) scales for a sample of six nearby barred galaxies. We collected multi-wavelength photometric data from far-ultraviolet (FUV) to far-infrared (FIR) from the DustPedia database and applied a panchromatic spectral energy distribution (SED) fitting procedure on square apertures of fixed angular size ( ) using the magphys code.
Results. For each galaxy, we obtain the distributions of stellar mass and SFR surface density and relate them in the – plane, deriving the spatially resolved MS relation. Although significant galaxy-to-galaxy variations are in place, we infer the presence of a common anti-correlation track in correspondence with the bar-hosting region, which shows systematically lower SFRs. This central quiescent signature can be interpreted as the result of a bar-driven depletion of gas reservoirs and a consequent halting of star formation. Our findings appear to support an inside-out quenching scenario.
@@
Internal processes typically dominate the late-stage evolution of galaxies, when gravitational interactions and mergers become less frequent as the Universe expands (Kormendy & Kennicutt, 2004). Stellar bars are among the main internal drivers of the secular evolution of disc galaxies, because they are able to shape the current properties of their hosts, redistributing the gas, stars, and angular momentum within them (e.g. Athanassoula & Misiriotis, 2002; Sheth et al., 2005; Debattista et al., 2006; Athanassoula, 2013; Di Matteo et al., 2013; Sellwood, 2014; Romeo et al., 2023).
@@
The pivotal role of bars in improving our understanding of galaxy evolution is also demonstrated by the fact that they are common morphological structures in the local Universe. According to optical and near-infrared (NIR) surveys, between and of nearby disc galaxies host stellar bars (e.g. Eskridge et al., 2000; Barazza et al., 2008; Aguerri et al., 2009; Nair & Abraham, 2010; Buta et al., 2015; Erwin, 2018). The evolution of this fraction over cosmic time may shed light on the epoch of bar formation. In particular, as found by several observational studies, there seems to be a consensus on a decline in the barred galaxy fraction from the local Universe up to by at least a factor of two (e.g. Sheth et al., 2008; Melvin et al., 2014). More recently, the unprecedented sensitivity and spatial resolution in the NIR of the James Webb Space Telescope (JWST) allowed authors to infer the evolution of the bar fraction at higher redshifts. Prominent elongated structures have been identified in disc galaxies up to (e.g. Guo et al., 2023; Costantin et al., 2023; Amvrosiadis et al., 2024), while Le Conte et al. (2023) presented a full census of observable stellar bars at . These latter authors pointed out that the incidence of bars in disc galaxies decreases from to going from to . Consistent results were also obtained in a fully cosmological context thanks to a new generation of high-resolution simulations (e.g. Fragkoudi et al., 2020; Zhao et al., 2020; Rosas-Guevara et al., 2022).
@@
The fact that both observational studies and simulations present evidence of a significant population of barred galaxies up to suggests that stellar bars are already in place in the early Universe, highlighting the possibility that bar-driven secular processes influence galaxy evolution over billions of years. Bars are thought to form over timescales of the order of a hundred million years, once a dynamically cold and rotationally supported disc has settled (e.g. Ostriker & Peebles, 1973; Sellwood & Wilkinson, 1993). The onset of bar formation occurs mainly when the stellar orbits of massive disc galaxies deviate from a circular path due to instabilities.
@@
Whether the growth of stellar bars enhances or suppresses star formation is still debated (e.g. Kim et al., 2017; Wang et al., 2020). However, there are indications supporting the hypothesis that bars can suppress star formation (bar quenching) in a region spanning from the central subkiloparsec (subkpc) to the end of the bar itself (e.g. Masters et al., 2010, 2012; Gavazzi et al., 2015; James & Percival, 2016; George et al., 2019a; Rosas-Guevara et al., 2020; Zhang et al., 2021), thus contributing to the observed transition from active star-forming galaxies to passive ones (e.g. Man & Belli, 2018). This could be due to the bar-induced shock and shear; these stabilise the gas against collapse by increasing turbulence, which, in turn, results in the inhibition of star formation (e.g. Haywood et al., 2016; Khoperskov et al., 2018). Another possible mechanism is related to the redistribution of cold gas along the length of the stellar bar. The torque produced by this morphological structure perturbs the interstellar medium (ISM), driving gas and dust inflow towards the galactic centre (e.g. Casasola et al., 2011; Combes et al., 2013; Spinoso et al., 2017). As a consequence, this boosts the star formation in the inner subkpc-scale area and depletes the bar-hosting region of fuel for further star formation (e.g. George et al., 2019a, 2020; Newnham et al., 2020).
@@
The bar-quenching scenario suggested by many recent observations and simulations predicts that bars can be responsible for the onset of a phase of high star formation rate (SFR) at their early evolutionary stages, potentially triggering starburst events or fuelling an active galactic nucleus (AGN), accompanied by a subsequent rapid consumption of gas, thus inducing a premature quenching phase in their host galaxies (e.g. Spinoso et al., 2017; Fraser-McKelvie et al., 2020a, b; George & Subramanian, 2021; Géron et al., 2021). This could explain why some literature works agree on the fact that barred galaxies have higher SFRs than unbarred ones, especially at the centre, along the length of the bar, and within the circumnuclear region (e.g. Sheth et al., 2005; Zurita & Pérez, 2008; Ellison et al., 2011; Coelho & Gadotti, 2011; Prieto et al., 2019; Lin et al., 2020; Sánchez-García et al., 2023). Conversely, many other studies have pointed out that bars are most likely to be found in quiescent galaxies, which are typically more massive, redder, and gas-poor (e.g. Vera et al., 2016; Cervantes Sodi, 2017; Géron et al., 2021).
@@
In this context, it is important to further explore the impact of bar quenching by studying the star-formation properties of barred galaxies and their dependence on the presence of the bar, especially through the existing empirical correlations between the main structural parameters of galaxies. In particular, the stellar mass ()–SFR relation is one of the most useful tools in modern astrophysics (e.g. Brinchmann et al., 2004; Noeske et al., 2007a, b; Salim et al., 2007). Habitually referred to as the main sequence (MS) of star-forming galaxies (SFGs), it shows that more massive galaxies host larger amounts of star formation than those of lower mass. Many literature works suggest that this log-linear correlation is valid —with an intrinsic scatter of 0.3 dex— for both local and distant galaxies (up to z 6) and for a wide range of stellar masses and considering different SFR tracers (e.g. Daddi et al., 2007; Pannella et al., 2009; Santini et al., 2009; Oliver et al., 2010; Elbaz et al., 2011; Lee et al., 2012; Whitaker et al., 2012; Moustakas et al., 2013; Whitaker et al., 2014; Tasca et al., 2015; Barro et al., 2017; Santini et al., 2017; Pearson et al., 2018; Popesso et al., 2019b; Daddi et al., 2022; Popesso et al., 2023). Galaxies exhibiting various rates of stellar production (starbursts, SFGs, and passive galaxies) populate different regions of the –SFR plane. Therefore, their position on the MS can be used as a diagnostic tool to identify their state of star formation and to figure out the origin of their colour bimodality (e.g. Rodighiero et al., 2011; Renzini & Peng, 2015; Bisigello et al., 2018; Popesso et al., 2019a).
@@
A significant limitation in our understanding of the –SFR relation stems from the fact that many studies have relied on the integrated properties of galaxies, without distinguishing among their morphological components. A more advanced approach involves characterising this correlation on smaller scales employing spatially resolved data. Thanks to high-spatial-resolution imaging and the advent of integral field spectroscopy (IFS) facilities, the MS relation was found to also be valid locally, at kpc and subkpc scales (e.g. Sánchez et al., 2013; Wuyts et al., 2013; Cano-Díaz et al., 2016; Abdurro’uf & Akiyama, 2017; Hsieh et al., 2017; Lin et al., 2019; Enia et al., 2020; Ellison et al., 2021; Casasola et al., 2022). This implies that the star-formation activity originates from physical mechanisms that could be universal across various physical scales. There is a general consensus on the existence of such a ‘resolved’ MS relation between the locally measured SFR surface density () and the stellar mass surface density (). However, the slope, zero-point, and scatter of this relation are subject to significant galaxy-to-galaxy variation, and depend on the different morphological structures, SFR tracers, dust-extinction correction, and fitting procedures (e.g. Abdurro’uf & Akiyama, 2017; Vulcani et al., 2019; Pessa et al., 2021, 2022; Casasola et al., 2022).
@@
With the aim of probing the relevance of bar-driven quenching, several observational works and simulations analysed the integrated MS relation of statistically large samples of barred galaxies (e.g. Vera et al., 2016; George et al., 2019b; Fraser-McKelvie et al., 2020a; George & Subramanian, 2021; Rosas-Guevara et al., 2020, 2022). However, to the best of our knowledge, there are just a few studies that investigated the spatially resolved MS across different galactic structures, including the bar (Pessa et al., 2021, 2022). Moreover, previous literature works are almost all based on spectroscopic data or UV-to-optical tracers of SFR (see Sánchez, 2020, for a review).
@@
In the present work, we further investigated the MS relation at subkpc scales for six barred galaxies in the local Universe, exploiting the power of multi-wavelength photometric data —from far-ultraviolet (FUV) to far-infrared (FIR)— collected in the DustPedia archive (Davies et al., 2017; Clark et al., 2018) and building on the methodology proposed by Enia et al. (2020). With this approach, we retrieve a reliable measure of the SFR, taking into account both the unobscured and the dust-obscured components, once a spectral energy distribution (SED) fitting procedure is applied. The final aim of this study is to get a more comprehensive view of the star-formation processes in barred galaxies and to understand whether the presence of the bar affects the typical log-linear trend of the resolved MS, at least in the inner regions.
@@
The paper is structured as follows: in Section 2, we briefly describe the data used in this work and our sample selection; in Section 3, we report the main steps of the methodology used to derive and SFR; in Section 4, we present our results concerning the spatially resolved MS; in Section 5, we discuss the implications of our findings. Finally, in Section 6, we summarise our work and the main outcomes. Throughout the paper, we assume a CDM cosmology with parameters , , and (Planck Collaboration et al., 2016), and a Chabrier (2003) initial mass function (IMF).
@@ 2.1 DustPedia archive
DustPedia111The DustPedia website is available at http://dustpedia.astro.noa.gr/ is a research project that was developed to enable a comprehensive characterisation of the cosmic dust in the local Universe and of its effects on multi-wavelength observations of ISM in galaxies (Davies et al., 2017). The DustPedia sample consists of 875 galaxies that were observed by Herschel (Pilbratt et al., 2010) and that lie within 50 Mpc, with an optical diameter D25222D25 is the major axis of the 25-th level isophote, which is the one at which the value of the surface brightness falls below 25mag/arcsec2 in the B band. of greater than . The complete data set, which is publicly available to the scientific community, contains standardised multi-wavelength imagery and photometry of the whole sample of galaxies, including both custom reductions and archival data, spanning over five orders of magnitude in wavelength from UV to microwave (Clark et al., 2018).
@@ 2.2 Sample selection
To properly select a sample of barred spiral galaxies from the DustPedia database, we first put some constraints on the Hubble stage index T444The Hubble stage index T is a numerical value, running from -6 to 10, that is assigned to each morphological class of galaxies in the De Vaucouleurs scheme. Negative values correspond to early-type galaxies (ellipticals and lenticulars), while positive numbers correspond to late-type galaxies (spirals and irregulars)., restricting its range between 2 and 7. This allows us to exclude early-type galaxies, intermediate classifications between lenticular and spirals, and also irregular galaxies. We then applied a cutoff in distance, at approximately 20 Mpc, and also imposed a lower limit on the optical diameter D25, considering only galaxies with D . Both these conditions ensure a selection of sufficiently close and extended galaxies, so that their central regions hosting the bar can be properly investigated with sufficient resolution. Furthermore, we focus on sources that have a disc inclination of 65°to avoid projection corrections. In addition, we selected only spiral galaxies that show a bar structure based on their known morphological classification. We also exclude objects with peculiar structures and those hosting obvious AGN (according to the literature). Indeed, the presence of strong nuclear activity would largely affect the emission from star formation in most of the galaxy and especially in the central region, which is the area we are investigating in order to probe the relevance of the bar-driven quenching. Finally, following these outlined constraints, we obtained a sample that is as free as possible from any contamination, enabling us to more effectively distinguish the impact of the bar.
@@
The final sample consists of six objects: NGC 1566, NGC 3351, NGC 3953, NGC 4579, NGC 4725, and NGC 5236. Distances, inclinations, sizes, and integrated properties of the galaxies under consideration are presented below and summarised in Table 1. All the sources are identified by their own equatorial coordinates in right ascension (RA) and declination (Dec) at epoch J2000. For each galaxy, along with the morphological type, we report: its luminosity distance in megaparsecs (Mpc), inclination, R25 (semi-major axis of the 25th level isophote), and size in kpc. All these quantities are taken from the HyperLEDA database (Makarov et al., 2014) and are those adopted by the DustPedia collaboration. Moreover, we report the bar radius Rbar in kpc, as determined in previous studies based on photometric decompositions of images from the Spitzer Survey of Stellar Structure in Galaxies (S4G). In particular, we retrieve the bar radius of NGC 1566 and NGC 3351 from Muñoz-Mateos et al. (2013), while we refer to the work of Salo et al. (2015) for details of the bars of the other galaxies. We then show the total stellar mass and the total SFR —with the corresponding uncertainties— of all the selected sources. These integrated properties are expressed in units of and yr-1, respectively, and are taken from the results of Nersesian et al. (2019), paying attention to properly rescale them in accordance with a Chabrier (2003) IMF. Nersesian et al. (2019) derived the global properties of most of the DustPedia sample using the aperture-matched photometry generated by the Comprehensive and Adaptable Aperture Photometry Routine (CAAPR), which is presented in Clark et al. (2018). On this photometric dataset, Nersesian et al. (2019) modelled the panchromatic SED employing the cigale fitting code (Boquien et al., 2019) and assuming the Salpeter (1955) IMF. We also indicate the size (in kpc) of the square apertures adopted for the spatially resolved analysis presented in this paper (i.e. the size of the cells). In particular, this value is related to the so-called pixel-by-pixel SED-fitting procedure that is applied to a grid of square cells of per side, which corresponds to different physical sizes depending on the distance of the galaxy. Further details of the methodology are provided in Sect. 3.